# Diffusion Model in Hyperspectral Image Processing and Analysis: A Review

### v1

@@ III-B Evolution of the Diffusion model with different variants
III-C Traditional Deep Learning Models
@@
Generative Models (GM) is a class of machine learning models that can learn distributions from data and generate new samples. It includes data generation, data enhancement, image restoration, anomaly detection and so on. It is mainly valuable in computer vision, natural language processing and other fields. In this paper, we will focus on the traditional generative models: the GAN, VAE, autoregressive models.
@@
Generative adversarial network (GAN) is another important network framework for unsupervised anomaly detection techniques.GAN was originally conceptualized in the zero-sum game framework of game theory. It was proposed by Ian Goodfellow et al. in 2014 [8]. The core idea is to generate data by adversarial training of two neural networks: a generator and a discriminator. The generator is responsible for generating data, and the discriminator is responsible for distinguishing the generated data from the real data. Through this adversarial process, the generator gradually learns to generate data similar to the distribution of the real data, which is fast to generate, but the training is unstable and prone to pattern collapse problem [58, 59, 60].
@@
Variational Auto-Encoder (VAE) [9] is a generative model that combines the advantages of Autoencoder and probabilistic graphical models to learn latent representations from data and generate new data samples.The core idea of VAE is to approximate the latent distributions of the data by Variational Inference (VII), and to generate the decoder by generating new data with lower generation quality but stable training [61, 62, 63].
@@
Autoregressive Models (ARMs) [64] are a class of generative models that model sequence data by assuming that the current data point is dependent on previous data points. The core idea of autoregressive models is to utilize the temporal dependence of sequence data to generate new data points step by step, e.g., RNN (Recurrent Neural Network) forms a recursive body in the process of processing the data; LSTM (Long Short-Term Memory Neural Network)) a kind of RNN with a special pattern. Autoregressive models generate the complete data by generating each part of the data step by step, which is slow in generating speed, but generating high quality [65, 66].
III-D Successful Application of Diffusion Modeling in Image Generation
Diffusion model has a wide range of applications, for images on the application began in 2020, mainly based on Denoising Diffusion Probabilistic Models (DDPM), first of all, the initial proposed DDPM [19], IDDM [44], and DDIM [35], etc., through the Markov chain as well as the addition of the noise and denoising process, and finally the prediction of the noise residuals to optimize the quality of the generation. However, the number of sampling steps is numerous and the computation speed is slow. Then ADM-G (Classifier-Guided-Diffusion) [11], introduced by OpenAI team in 2021, made the model surpass GAN in FID metrics for the first time by introducing classifier-guided generation. but dependent on classifiers and high computational cost. Later, in order to optimize text-image alignment and super-resolution generation, Imagen [46] was introduced by the Google Research team, which combines the T5-XXL text encoder with the cascade diffusion architecture, and has outstanding text alignment capability, but with great resource requirements. Later, in order to realize multimodality and efficient generation, Stabel-Diffusion [18] was proposed by Robin Rombach et al. in 2023, which compressed the diffusion process into the latent space, significantly reducing the computational complexity, but losing information in the latent space. Meanwhile, also by the OpenAI team, CLIP-Guded-Diffusion [49] was introduced, which combines the CLIP model to realize cross-modal generation, supports complex semantic control, and generates semantic consistency, but has large training requirements and slow generation speed. Subsequently, in order to unify the framework and extend the multi-tasking domain, OmniGen [53], the first unified framework supporting tasks such as text generation, image editing, and so on, without additional plug-ins, was introduced by the Wisdom Source Institute in 2024. Then, the Byte Jump team introduced D-Dit [55], which combines continuous and discrete diffusion techniques to achieve bidirectional text-image generation, support multimodal tasks, and improve parameter efficiency by about 20%. In 2025 by Yifan Zhou et al. proposed AF-LDM [57], which introduces shift isovariance loss, improves the UNet attention module, reduces potential spatial noise bias, improves video generation consistency, and enhances the model robustness, etc.
At the same time, the diffusion model will continue to be innovated and developed in the hyperspectral field in the future, and these models are improved on the basis of the main diffusion model, aiming at solving the problem of image in classification, anomaly detection and recovery, and so on.
@@ III-D Successful Application of Diffusion Modeling in Image Generation
TABLE I: Basic Diffusion Models and Variants
@@
IV Applications in Hyperspectral Image Processing and Analysis
Deep learning models are widely used in various fields and show powerful advantages, in recent years, some scholars have also applied them to the field of hyperspectral, while making many achievements. The application in hyperspectral image analysis is embodied in many aspects, mainly including data enhancement and synthesis, generation and complementation of hyperspectral data, image classification, segmentation, target detection, and anomaly detection, and many other aspects, in this section, the hyperspectral image processing is divided into two aspects: one is the image processing under the traditional deep learning model, and the other is the image processing under the diffusion-based model. The following is a description of these methods.
@@ IV Applications in Hyperspectral Image Processing and Analysis
IV-A The main Diffusion model used in the field of hyperspectral imagery
Diffusion Model based Hyperspectral Image Processing, which was first applied in the field of Hyperspectral in 2022, meanwhile, it was obtained through the research that the main use of Diffusion Modeling methods in this field is mainly Denoising Diffusion Probabilistic Models (DDPM), and this paper also presents the technical analysis of DDPM and its variants in the field in this section. The main diffusion models for hyperspectral image processing based on diffusion modeling are DDPM [19], IDDPM [44], DDIM [35], LDM [18], Classifier-Gided [11], Stable-Diffusion [18].
First of all, the basic principles of DDPM have been introduced above, and the basic principles and formulas of IDDPM and DDIM are the same as those of DDPM, so we won’t repeat them here.LDM and Stable-Diffusion belong to the same category, in which the core ideas are the same, and it can be said that Stable-Diffusion is the specific implementation version of LDM, so the following will introduce the following LDM (Latent Diffusion Model), Classifier-Guided (Conditional Diffusion Model) two basic principles, formula.
IV-A1 Conditional Diffusion Model (CDM)
Conditional diffusion models are similar to the development of GANs in that diffusion models begin with unconditional generation, followed immediately by conditional generation [67]. Unconditional generation is often used to explore the upper limits of the model’s capabilities, whereas conditional generation is more beneficial for applications as it allows the output to be controlled according to the human’s wishes.
The first work to introduce this condition in diffusion models was [11], which guided the generation of diffusion models by adding classifiers to trained diffusion models, hence the term guided diffusion models. Although this approach has a low training cost, it increases the inference cost by using the classification results to guide the sampling process of the diffusion model and has poor control over the details to produce images with better results. Therefore, the Google team [68] decided to adopt a simple idea of controlling the generated results by conditionally retraining the DDPM and named it Classifier-Free Guidance.
Formally, given the conditional information $\mathbf{c}$, the distribution of the combined DDPM changes from equations (2)(3), and accordingly, the optimization objective and sampling process are modified to the following form:
@@ IV-A1 Conditional Diffusion Model (CDM)
This approach is more widely used than guided diffusion models and is the basis for numerous models of important diffusion (e.g., DALL-E2 [45], Imagen [49], and stabilizing diffusion (SD) [18]).
IV-A2 Latent Diffusion Model (LDM)
The computational cost of direct manipulation of the diffusion model in the original image pixel space is high, Rombach et al [60] proposed to train and reason about the diffusion model in a low-dimensional latent space and named it Latent Diffusion Model (LDM).
Figure 7: Architecture of LDM, where E and D denote the encoder and decoder, respectively
@@ IV-A2 Latent Diffusion Model (LDM)
As shown in Fig. 7, the LDM consists of three parts. First, the specific diffusion process of the LDM, which shifts the workspace from pixel space to latent space using a pre-trained autoencoder, is shown in the leftmost gray box in Fig. 7, where the noise is added to feature $\mathbf{z}$ to get the lower half of $\mathbf{z_{T}}$ as forward diffusion; the middle orange box is in latent space, where the U-Net with crosscutting attention supports the multimodal inputs in order to restore $\mathbf{z_{T}}$ to $\mathbf{z}$ as reverse diffusion. The right side box is the conditional encoder, which encodes various conditions (e.g., text and image) into feature vectors $\tau_{\theta}$ and feeds them into U-Net to generate an image.LDM is able to effectively reduce computational cost without degrading the quality of image synthesis while speeding up the inference speed, which is widely used in natural scenes and provides inspiration for the application of diffusion models to hyperspectral applications.
IV-B Application of Diffusion model in image processing
IV-B1 Data Enhancement and Synthesis
@@
In data enhancement and synthesis, usually super-resolution techniques have an important role and unique advantages in improving quality. Super-resolution aims at reconstructing high-resolution images from low-resolution images [69], but the images can be disturbed, due to imaging equipment, weather conditions or downsampling. The loss of detail in hyperspectral images can be significant when compared to natural images, making inverted reconstruction of high-resolution images more challenging. Before the advent of diffusion models, deep neural network based approaches, e.g., Convolutional Neural Networks (CNNs) and Transformers, were finding suitable mapping functions to integrate feature information into a single step. To achieve this goal, methods including self-supervised learning [70], adaptive marker selection [71] and frequency information assistance [72] were used. However, the diffusion model-based approach fuses low-resolution images and bootstrap images in each step of the diffusion process, which helps to better integrate different image information and, at the same time, has gained wide attention in the hyperspectral field.
Although HSI has a high spectral resolution, its spatial resolution is relatively low [73], which may limit the performance of various HSI-based applications. To obtain high-resolution hyperspectral maps, there are two types of methods: panchromatic sharpening [74] and multispectral and HSI fusion [75]. The former enhances HSI by injecting detail information in the panchromatic (PAN) image, while the latter utilizes multispectral images to help hyperspectral maps learn spatial details. For example, Shi et al [76] used multispectral and hyperspectral images in tandem as conditional inputs to a diffusion model, and then generated hyperspectral maps with high resolution from the information found in the two image modalities.Qu et al [77] devised a conditional recurrent diffusion framework that allows the spatial SR process and the spectral SR process to be guided by the multispectral image and the hyperspectral map, respectively, to generate high-resolution hyperspectral maps with complementary spatial and spectral information to generate high-resolution hyperspectral maps.
@@
More contemporary research focuses on the use of PAN image HSI super-resolution [78, 79, 80, 81, 82, 83, 84].Meng et al [79] proposed a modal calibration module to enhance and extract features in two images, where the enhanced features are used as conditional inputs to the diffusion model.Cao et al [78] argued that the different image modal unique information should not be mixed for processing and designed two conditional modulation modules to extract coarse-grained style information and fine-grained frequency information as conditional inputs.
@@
In order to take full advantage of the unique information of different modalities, Li et al [82] proposed a panchromatic sharpening network based on a biconditional diffusion model, where HSI and PAN are used as independent conditional inputs to learn the spectral features and spatial texture, respectively. The network structure is designed based on LDM and achieves excellent SR performance while significantly reducing the computational cost.Rui et al [84] also perform the sampling process in low-dimensional space. Based on the fact that the high-resolution hyperspectral map can be decomposed into the product of two low-rank tensors, one of the low-rank tensors is first computed using the low-resolution hyperspectral map. This tensor, together with the LRHSI and PAN, is then conditioned and fed into a pre-trained diffusion model [85] for the other low-rank tensor. Unlike the above methods, this method is a completely unsupervised deep learning method that does not require the involvement of high-resolution hyperspectral maps in any of the processes, providing feasibility for its application in practice.
IV-B2 Data Generation and Completion
Generation and Complementation in general can be categorized as text-to-image and image-to-image, but in the hyperspectral domain only image-to-image exists. For diffusion models, diffusion models have image bootstrapping to generate new images. Guided images can be in various forms such as maps [86], semantic layouts [87, 88, 89, 90]. Even though these images contain only some specific information, they still show excellent performance in image generation. For example, Espinosa and Crowley [86] proposed to use maps and historical maps to train conditional diffusion models, which generated more realistic satellite images.Yuan et al. [87] generated high-quality hyperspectral images under the guidance of semantic masks, and also solved the intrinsic problem that diffusion models require a long time of training for model convergence. However, these things alone are not enough, so it is important to generate not only high-quality images, but also image-annotation pairs that are more practical. For example, Zhao et al [89] implemented a two-stage fine-tuning of the SD model to enable the generation from noise to annotations and then from annotations to images.Toker et al [90] extended the standard diffusion model to a joint probabilistic model for images and their corresponding labels, which enabled the simultaneous generation of labels and images.

### v2

@@ 3.2 Evolution of the Diffusion model with different variants
3.3 Traditional Deep Learning Models
@@
Generative Models (GMs) are a class of machine learning models capable of learning data distributions and generating novel samples. Their applications are widespread in fields like computer vision and natural language processing, encompassing tasks such as data generation, data enhancement, image restoration, and anomaly detection. In this paper, we will primarily focus on traditional generative models, specifically Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and autoregressive models.
@@
The Generative Adversarial Network (GAN), proposed by Ian Goodfellow et al. in 2014 [8], is another significant framework for unsupervised anomaly detection, rooted in the concept of a zero-sum game. At its core, a GAN generates data through the adversarial training of two neural networks: a generator and a discriminator. The generator’s role is to produce data, while the discriminator’s task is to differentiate between this generated data and real data. This adversarial process drives the generator to progressively learn how to create data that closely resembles the distribution of real data. While GANs offer fast generation, they often suffer from unstable training and are prone to issues like mode collapse [58, 59, 60].
@@
The Variational Auto-Encoder (VAE) [9] is a generative model that blends the strengths of Autoencoders with probabilistic graphical models. Its core idea is to learn latent representations from data and subsequently generate new samples. VAEs achieve this by using Variational Inference (VI) to approximate the data’s latent distributions. While they offer stable training, the quality of the generated data can sometimes be lower [61, 62, 63].
@@
Autoregressive Models (ARMs) [64] are a class of generative models that operate by assuming the current data point in a sequence is dependent on preceding ones. Their core idea is to leverage this temporal dependence to generate new data points sequentially. Examples include Recurrent Neural Networks (RNNs), which employ a recursive structure for data processing, and Long Short-Term Memory (LSTM) Neural Networks, a specialized type of RNN. While ARMs are known for generating high-quality data, their step-by-step generation process can result in slower speeds [65, 66].
3.4 Successful Application of Diffusion Modeling in Image Generation
Since 2020, Diffusion Models have seen widespread application in image processing, primarily leveraging Denoising Diffusion Probabilistic Models (DDPMs). Early works like the initial DDPM [19], IDDM [44], and DDIM [35] optimized generation quality through Markov chains and refined noise addition/denoising processes, predicting noise residuals. However, these models were characterized by numerous sampling steps and slow computation.
In 2021, the OpenAI team introduced ADM-G (Classifier-Guided-Diffusion) [11], which, by integrating classifier-guided generation, allowed the model to surpass GANs in FID metrics for the first time. Despite this breakthrough, it remained dependent on classifiers and incurred high computational costs. To enhance text-image alignment and super-resolution, the Google Research team later introduced Imagen [46], combining the T5-XXL text encoder with a cascaded diffusion architecture, resulting in outstanding text alignment capabilities but demanding substantial resources.
Seeking multimodal and efficient generation, Robin Rombach et al. proposed Stable Diffusion in 2023 [18]. This model significantly reduced computational complexity by compressing the diffusion process into a latent space, albeit with some information loss. Concurrently, the OpenAI team also released CLIP-Guided-Diffusion [49], which integrates the CLIP model to enable cross-modal generation and support complex semantic control, ensuring semantic consistency. However, this model also requires significant training resources and has a slower generation speed.
More recently, in 2024, the Wisdom Source Institute introduced OmniGen [53], the first unified framework supporting tasks like text generation and image editing without additional plug-ins. Following this, the ByteDance team launched D-Dit [55], which combines continuous and discrete diffusion techniques to achieve bidirectional text-image generation, support multimodal tasks, and improve parameter efficiency by approximately 20%. Looking ahead to 2025, Yifan Zhou et al. are set to propose AF-LDM [57], aiming to enhance video generation consistency and model robustness by introducing shift isovariance loss and improving the UNet attention module to reduce potential spatial noise bias.
Simultaneously, the diffusion model is poised for continued innovation and development within the hyperspectral domain. Future advancements in this area will build upon existing diffusion models, with a focus on resolving challenges in hyperspectral image classification, anomaly detection, and recovery.
@@ 3.4 Successful Application of Diffusion Modeling in Image Generation
Table 1: Summary of basic diffusion models, variants and their characteristics
@@
4 Applications in Hyperspectral Image Processing and Analysis
Deep learning models are widely adopted across various fields, demonstrating significant advantages. In recent years, scholars have successfully applied them to hyperspectral imaging, yielding numerous achievements. Their application in hyperspectral image analysis spans several areas, including data enhancement and synthesis, data generation and completion, image classification, segmentation, target detection, and anomaly detection. This section will categorize hyperspectral image processing into two main approaches: those using traditional deep learning models and those employing diffusion-based models. The following provides a detailed description of these methods.
@@ 4 Applications in Hyperspectral Image Processing and Analysis
4.1 The main Diffusion model used in the field of hyperspectral imagery
Diffusion Model based Hyperspectral Image Processing was first applied to the hyperspectral field in 2022. Research indicates that Denoising Diffusion Probabilistic Models (DDPM) are predominantly used in this domain. This section will specifically analyze the technical aspects of DDPM and its variants within hyperspectral image processing. The primary diffusion models leveraged for this purpose include DDPM [19], IDDPM [44], DDIM [35], LDM [18], Classifier-Guided Diffusion [11], and Stable Diffusion [18].
We’ve already covered the basic principles of DDPMs. Since the fundamental principles and formulas of IDDPM and DDIM are essentially identical to DDPM, we won’t reiterate them here. LDM (Latent Diffusion Model) and Stable-Diffusion fall under the same category, sharing core ideas, with Stable-Diffusion effectively being a specific implementation of LDM. Therefore, the following sections will introduce the basic principles and formulas of LDM and Classifier-Guided (Conditional Diffusion Model).
4.1.1 Conditional Diffusion Model (CDM)
Similar to the evolution of GANs, diffusion models initially focused on unconditional generation, swiftly transitioning to conditional generation [67]. While unconditional generation is often used to push the boundaries of a model’s capabilities, conditional generation holds more practical value for applications, as it enables the output to be controlled based on human intent.
The introduction of conditioning into diffusion models first appeared in [11]. This pioneering work guided the generation of diffusion models by incorporating classifiers into pre-trained models, leading to the term "guided diffusion models." While this method offered low training costs, it increased inference costs by using classification results to steer the diffusion model’s sampling process, and it struggled to precisely control details for optimal image output. Consequently, the Google team [68] opted for a simpler approach: retraining the DDPM conditionally to control generated results, a technique they named Classifier-Free Guidance.
Formally, when given conditional information $\mathbf{c}$, the combined DDPM distribution changes from equations (2) and (3). Consequently, both the optimization objective and the sampling process are modified to take the following form:
@@ 4.1.1 Conditional Diffusion Model (CDM)
This approach has seen broader adoption than guided diffusion models and forms the foundation for many significant diffusion models, including DALL-E2 [45], Imagen [49], and Stable Diffusion (SD) [18].
4.1.2 Latent Diffusion Model (LDM)
Directly manipulating diffusion models in the original image pixel space incurs high computational costs. To address this, Rombach et al. [60] proposed training and performing inference with diffusion models within a low-dimensional latent space, terming this approach the Latent Diffusion Model (LDM).
Figure 7: Architecture of LDM, where E and D denote the encoder and decoder, respectively [18]
@@ 4.1.2 Latent Diffusion Model (LDM)
As depicted in Fig. 7, the Latent Diffusion Model (LDM) comprises three distinct parts. Firstly, the leftmost gray box illustrates the LDM’s specific diffusion process: it shifts its operational domain from pixel space to a low-dimensional latent space by utilizing a pre-trained autoencoder. Within this space, noise is incrementally added to feature $\mathbf{z}$, yielding $\mathbf{z_{T}}$ as part of the forward diffusion. Secondly, the central orange box, also in latent space, houses a U-Net with cross-attention that supports multimodal inputs. This U-Net facilitates the reverse diffusion, restoring $\mathbf{z_{T}}$ back to $\mathbf{z}$. Lastly, the rightmost box contains the conditional encoder, which transforms various conditions (such as text and images) into feature vectors $\tau_{\theta}$. These vectors are then fed into the U-Net to guide image generation. The LDM’s ability to effectively reduce computational cost without sacrificing image synthesis quality, coupled with its accelerated inference speed, has led to its widespread adoption in natural scene generation and offers significant inspiration for applying diffusion models to hyperspectral applications.
4.2 Application of Diffusion model in image processing
4.2.1 Data Enhancement and Synthesis
In data enhancement and synthesis, super-resolution techniques play a crucial role in improving image quality. Super-resolution aims to reconstruct high-resolution images from their low-resolution counterparts [69]. However, these images can be degraded by factors like imaging equipment, weather conditions, or downsampling. This loss of detail is often more pronounced in hyperspectral images compared to natural images, making the inverse problem of reconstructing high-resolution versions significantly more challenging.
@@ 4.2.1 Data Enhancement and Synthesis
Before the advent of diffusion models, approaches based on deep neural networks, such as Convolutional Neural Networks (CNNs) and Transformers, sought to integrate feature information in a single step by finding suitable mapping functions. Methods like self-supervised learning [70], adaptive marker selection [71], and frequency information assistance [72] were employed to achieve this. In contrast, diffusion model-based approaches fuse low-resolution images and bootstrap images at each step of the diffusion process. This iterative integration of different image information has proven highly effective and has garnered significant attention in the hyperspectral field.
Although Hyperspectral Images (HSIs) boast high spectral resolution, their spatial resolution is often comparatively low [73], which can limit the effectiveness of various HSI-based applications. To overcome this, two primary methods are employed to achieve high-resolution hyperspectral maps: panchromatic sharpening [74] and multispectral and HSI fusion [75].
Panchromatic sharpening enhances HSIs by injecting detailed spatial information from a high-resolution panchromatic (PAN) image. Conversely, multispectral and HSI fusion leverages multispectral images to help hyperspectral maps acquire spatial details. For instance, Shi et al. [76] utilized both multispectral and hyperspectral images as conditional inputs to a diffusion model, subsequently generating high-resolution hyperspectral maps by integrating information from these two image modalities. Similarly, Qu et al. [77] developed a conditional recurrent diffusion framework. This framework enables the spatial super-resolution process to be guided by the multispectral image and the spectral super-resolution process by the hyperspectral map, leading to the generation of high-resolution hyperspectral maps that boast complementary spatial and spectral information.
@@
More contemporary research is increasingly focused on utilizing Panchromatic (PAN) images for HSI super-resolution [78, 79, 80, 81, 82, 83, 84]. For instance, Meng et al. [79] introduced a modal calibration module to enhance and extract features from both image types, subsequently using these enhanced features as conditional inputs for a diffusion model. Conversely, Cao et al. [78] posited that the unique information from different image modalities should not be mixed during processing. They instead designed two conditional modulation modules to separately extract coarse-grained style information and fine-grained frequency information, which are then used as conditional inputs.
@@
To fully leverage the unique information from different modalities, Li et al. [82] proposed a panchromatic sharpening network based on a biconditional diffusion model. In this architecture, both HSI and PAN images serve as independent conditional inputs, allowing the network to separately learn spectral features and spatial textures. Designed based on LDM, this network achieves excellent super-resolution (SR) performance while significantly reducing computational cost. Similarly, Rui et al. [84] also perform the sampling process in a low-dimensional space. Their method is rooted in the fact that a high-resolution hyperspectral map can be decomposed into the product of two low-rank tensors. First, one of these low-rank tensors is computed using the low-resolution hyperspectral map. This tensor, along with the LRHSI (low-resolution hyperspectral image) and PAN, is then conditioned and fed into a pre-trained diffusion model [85] to generate the other low-rank tensor. Unlike the aforementioned methods, this approach is a completely unsupervised deep learning method, meaning it doesn’t require high-resolution hyperspectral maps during any part of the process, making it highly feasible for practical applications.
4.2.2 Data Generation and Completion
In the realm of generation and complementation, while broad categories include text-to-image and image-to-image synthesis, the hyperspectral domain exclusively features image-to-image applications. Here, diffusion models leverage image bootstrapping to create new images. These models can be guided by various forms of input, such as maps [86] or semantic layouts [87, 88, 89, 90]. Even with only specific information in these guiding images, excellent performance in image generation is still observed. For instance, Espinosa and Crowley [86] demonstrated that training conditional diffusion models with maps and historical maps could generate more realistic satellite images. Similarly, Yuan et al. [87] successfully produced high-quality hyperspectral images guided by semantic masks, simultaneously addressing the inherent challenge of lengthy training times for diffusion model convergence.
However, generating high-quality images alone isn’t always sufficient; creating more practical image-annotation pairs is often crucial. To this end, Zhao et al. [89] implemented a two-stage fine-tuning of the Stable Diffusion (SD) model, enabling generation from noise to annotations and then from annotations to images. Furthermore, Toker et al. [90] extended the standard diffusion model into a joint probabilistic model for images and their corresponding labels, allowing for the simultaneous generation of both.
Hyperspectral images (HSIs) offer rich spectral and spatial information, but their acquisition cost is significantly higher than that of multispectral images. Consequently, scientists have explored using readily available multispectral remote sensing images to generate HSIs [91, 92]. However, a challenge arises because the diffusion models used for this task require the input noise dimensions to match the spectral bands of the hyperspectral map. This can lead to an excessively large noise sampling space, making model convergence difficult.